please help me asap

According to a 2018 report from the Center on Education and the Workforce at Georgetown University, 70 percent of full-time college students are working. Suppose 15 students full-time college students are randomly selected and asked whether they are working. (a) (3 points) Explain (as best as you can) why this is a binomial experiment.
Answer #1

Solution:

Given that,

P = 0.70

1 - P = 0.30

n = 15

Here, BIN ( n , P ) that is , BIN (15 , 0.70)

then,

n*p = 15 * 0.70 = 10.5 > 5

n(1- P) = 15 * 0.30 = 4.5 < 5

Both conditions are not greater than 5, its follows binomial distribution.

Then this is a binomial experiment.